A 10.8 kg block is placed at rest on a plane inclined at 29 o. If after moving 6 meters, the block has a velocity of magnitude 6.3 m/s, what is the coefficient of friction between the block and the plane?
Answer #1

friction force f = uN

N = mgcos29  

so f = umgcos29

using work energy theorem,

work done by gravity + work done by friction = change in KE

mgLsin29   + ( - umgcos29 L ) = mv^2 /2 -   0

(9.81 x 6 x sin29) - ( u x 9.81 x cos29 x 6) = 6.3^2 /2   - 0

28.54 - u51.48 = 19.845

u = 0.17